Calculate the pka and ka given the pH.
Please show ALL steps
measured pH = cal, pKa = q. I 0 pH1 after 5 drops of NaOH-14- pH after 5 drops of HCl = gng BE SURE TO GET T 3) Buffer System with 2 ml acid + 20 ml conjugate base [B]/ [HB] = 10 BE SURE TO GET THIS CORRECT!! (use H+H Eq. above to calculate pKa) measured pH- 450 pH after addition of XS NaOH-10.00 dition of XS NaOH = calculated Ka - calculated. pKa = Show Calculation below leg 6,S
Answer #1

pH = 9.5

[B] / [HB] = 10 = [conjugate base / acid ]

pH = pKa + log [conjugate base / acid ]

9.50 = pKa + log 10

8.50 = pKa + 1

pKa = 8.50

pKa = -logKa

Ka = 10^-pKa

Ka = 10^-8.50

Ka = 3.16 x 10^-9

so

Ka = 3.16 x 10^-9

pKa = 8.50